HUBBLE SPACE TELESCOPE DAILY REPORT #2305 PERIOD COVERED: 0000Z (UTC) 02/09/99 - 0000Z (UTC) 02/10/99 Daily Status Report as of 041/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 6604 (Photometry of the Oldest Field Population in the Magellanic Clouds.) The WF/PC-2 was used to make both visual and inferred photometric observations of selected stars in the oldest fields in the Large Magellanic Cloud SMC 004600 and SMC 004612. From these data, the ages and population densities of the cluster will be measured. The observations were completed as scheduled, and no problems were noted. 1.5 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.6 Completed Two Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.7 Completed STIS/CCD/MA2 7400 (Evidence from Boron for Massive Star Evolution) The Space Telescope Imaging Telescope (CCD/MA2) was used to observe the B III and B II resonance lines at 2065.8 and 1362.5 Angstroms, respectively, in a sample of the main-sequence and slightly evolved blue star HD36591 using the HST STIS MAMA detectors and echelle gratings. A wave length calibration image was also taken. As described in 2.1 and HSTAR 6980, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the proposal completed with no further reported problems. 1.8 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. As described in 2.1 and HSTAR 6980, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.9 Completed WF/PC-2 6356 (SNAPSHOT Survey of Ultraluminous Infrared Galaxies) The WF/PC-2 was used to make observations of GAL-232201 as part of a survey of ultraluminous infrared galaxies. The nature of ultraluminous infrared galaxies has generated considerable controversy. The two main issues are: {1} are they powered by violent bursts of star formation or by hidden quasars? {2} are interactions and mergers invariably involved? The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2/STIS/CCD 8018 (Demographics of Nuclear Black Holes) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make observations to study the nuclear region of the normal galaxy NGC-4742 in order to study the demographics of black holes. Detection of supermassive black holes in normal galaxies is one of the grand challenges that HST was designed to meet. A wavelength calibration observations was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 7929 (Cycle 7 WF/PC-2 CTE Monitor) The WF/PC-2 was used to make observations of NGC-5139 to make calibrations observations to monitor the CTE. This Cycle 7 proposal is executed every six months to monitor changes. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 The acquisition at 040/1140Z defaulted to fine lock backup on FGS-1 only when FGS-3 did not acquire its star. HSTAR 6980 was written. The proposals described in 1.7 and 1.8 may have been affected. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Operations Notes: The SSR EDAC error counter was reset once using ROP SR-1. A TTR was opened to document data inversion occurring on day 040. The STIS MCE-1 reset at 040/162812Z while in an SAA contour and while the low voltage was on. HSTAR 6979 was written. ROP NS-12 was executed to clear the error messages. MAMAM-1 will be recovered via SMS commanding at 041/1124Z. The temperature sensor for NICMOS FPA detector-2 was enabled at 040/1754Z. Per HSTAR 6981, the sensor on disabled at 040/2000Z when the temperature exceeded 300 degrees K. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH